Every story and comment from the Hacker News API, in a layout borrowed from 2001.
A comment on When intuition and math probably look wrong. Parent comment
The probability they are both heads is 1/2.
You're trying to formulate this as ....
P(head AND head | one is a head) = P(head AND head) / P(one is a head) = 1/4 * 4/3 = 4/12 = 1/3
P(one is a head) = 1 && P(head_a AND head_b) = P(head_a) = P(head_b)
There are no comments on this comment yet.
The probability they are both heads is 1/2.
You're trying to formulate this as ....
But this is wrong ... you know that you have a head ... which makes ... 1/3 would be the probability only on the first try (instead of stopping when you've got a head).